OpenStudy (anonymous):

6 + 10 x + 14 x^2 what's the GCF here and is it smaller than the smallest coefficient?

OpenStudy (amistre64):

2 see it as GCF = 2

OpenStudy (amistre64):

1,2,3,6 is the best you can get, what number do the rest have in common with this?

OpenStudy (anonymous):

I'm thinking

OpenStudy (amistre64):

....think....quieter.....

OpenStudy (anonymous):

It hurts....

OpenStudy (anonymous):

I am not sure...why do I not get this stuff?

OpenStudy (amistre64):

tell me, what are the coeeficients?

OpenStudy (anonymous):

2

OpenStudy (amistre64):

2 is a coefficient?

OpenStudy (anonymous):

I know whatever it is is not smaller then it is, because they are equal to each other...just not sure what number

OpenStudy (anonymous):

it is 6?

OpenStudy (amistre64):

6 is the smallest coefficient..yes; so any factors cannot possibly be larger than 6.

OpenStudy (anonymous):

Thank you so much..sorry about any trouble I caused

OpenStudy (amistre64):

think of it this way: if someone wants you to give them $10 and all you have is $3; what is the most you can give them?

OpenStudy (anonymous):

3

OpenStudy (amistre64):

exactly; so, in this problem of yours we have some numbers that are carrying cash... the largest amount of cash that they have in common cant be greater than the "$6" number right?

OpenStudy (anonymous):

right

OpenStudy (amistre64):

so lets open up that "$6" dollars wallet and see what hes got in there... hes got the following bills: $1, $2, $3 , $6. the rest of the numbers have similar types of cash. but the biggest numbered bill they have in common is a $2.

OpenStudy (anonymous):

I get it....that sounds simple...till you get to the way it looks in front, if they would word the things like you just did, I would be fine...

OpenStudy (amistre64):

the "10" is carrying: $1, $2, $5, $10 the "14" is carrying: $1, $2, $7, $14 the only "bills" they have in common are $2 bills

OpenStudy (anonymous):

Maybe you can help me with the next 2 things, do you have time?

OpenStudy (anonymous):

I wish you were my teacher.....

OpenStudy (amistre64):

i got at most an hour and a half :)

OpenStudy (anonymous):

only 2 small questions left....

OpenStudy (anonymous):

what are the different ways a quadratic polynomial can be factored out. What applies when?

OpenStudy (anonymous):

and.......... x ^ 3 + 2x^2 + x = 0 Can factorization still help solve this equation ? If yes, how ?

OpenStudy (anonymous):

Those are it...If you have time.

OpenStudy (amistre64):

lets dooooo this!!! ... :)

OpenStudy (amistre64):

thats my machoman randy savage impersonation :)

OpenStudy (amistre64):

what are the different ways a quadratic polynomial can be factored out. What applies when? the different ways to factor a quad equation..... well, there is the trial and error method

OpenStudy (anonymous):

I love your enthusiasm:)

OpenStudy (amistre64):

lol

OpenStudy (anonymous):

I think they should put a like button on here!lol

OpenStudy (amistre64):

there is the "completeing the square" method

OpenStudy (amistre64):

hah...i think thats the fan button. they give you cool titles like lifesaver, champion, star...but when you get to 100 and above, they turn you into a sandwich :/

OpenStudy (anonymous):

:( (to the sandwich part) I'm a fan of you:)

OpenStudy (anonymous):

I suck i have no fans

OpenStudy (amistre64):

awww.... thanx :) I hit 200 a few days ago and was hoping for something cool like "turkey club on rye"..but noooooo.....they stuck me with "superhero" lol

OpenStudy (anonymous):

be right back...I need to check on my daughters...

OpenStudy (anonymous):

lol...

OpenStudy (anonymous):

ok..they are fine

OpenStudy (amistre64):

the four most used methods to factor quadratics are: Factoring Factor by Grouping Quadratic Formula Completing the Square

OpenStudy (anonymous):

Does that have to do with the 1st question

OpenStudy (amistre64):

When do we use them? dunno really. Factoring is just splitting it up into its basic components... that can be done if its easy to recognize

OpenStudy (amistre64):

yeah sure...maybe :)

OpenStudy (amistre64):

by grouping generally ocuurs when there are more than 3 terms. we group them into pairs and factor them seperatley

OpenStudy (amistre64):

the quad formula is used when we just dont have time or patience to do the other ways :)

OpenStudy (amistre64):

and completing the square is just the long hand version of the quad formula, I guess you use it when you cant recall the quadratic formula :)

OpenStudy (anonymous):

Now to the 2nd question

OpenStudy (amistre64):

x ^ 3 + 2x^2 + x = 0 Can factorization still help solve this equation ? If yes, how ? Yes, if we can notice that all the terms include an "x" factor then we can remove it and do our quadratic techniques on the remaining terms like this: x ^ 3 + 2x^2 + x = 0 x(x^2 +2x +x) = 0 x (x+1)(x+1) = 0 x = 0 or x = -1

OpenStudy (amistre64):

that little straggling +x should be a +1....

OpenStudy (amistre64):

x(x^2 +2x +x) = 0 <- right in there x(x^2 +2x +1) = 0 x (x+1)(x+1) = 0

OpenStudy (anonymous):

ok, Now I am confused...

OpenStudy (anonymous):

lol...story of my life

OpenStudy (amistre64):

tell me whats got you confused :)

OpenStudy (anonymous):

about the straggling....

OpenStudy (amistre64):

I had a typo in there; that x^2 +2x +[x] <- right there. it should have been: x^2 +2x +1

OpenStudy (anonymous):

Thank you got it! I will probably be on again tomorrow.....I have this stuff all week and get really frustrated! Thanks for all your help! You are great! I am sure you are better then a sandwich:)

OpenStudy (amistre64):

thanx :) take care ;)

OpenStudy (anonymous):

You too